Problem 5. (4 pt) Find the general solution of y – 4y = 3e21.
Answer #1

y - uy =3024 crivon, Som d24 3e24 an2 - ug (8²y-uy) = 3 e 24 3e24. (02-4):y AE, 22-u=0 D - 12 cofa ciern - 2x + Ce - P.I 3e2:cheneral solution - cf + P.I Ciern 24 + cze + zne 8 4